OpenStudy (anonymous):

What is 3 root of 384? A. 3 3 root of 6 B. 4 3 root of 3 C. 4 3 root of 6 D. 6 3 root of 3

OpenStudy (anonymous):

it is supposed to look like this \[a.3\sqrt[3]{6}\]

OpenStudy (anonymous):

\[b.4\sqrt[3]{3}\]

OpenStudy (anonymous):

\[c.4\sqrt[3]{6}\]

OpenStudy (anonymous):

\[d.6\sqrt[3]{3}\]

OpenStudy (phi):

you can test each answer, by "moving" the number on the outside of the cube root to inside *but* put in 3 copies for example \[ 3\sqrt[3]{6} = \sqrt[3]{6\cdot 3\cdot 3\cdot 3 } \] is that the original problem? if we multiply it out we get \[ 3\sqrt[3]{6} = \sqrt[3]{6\cdot 3\cdot 3\cdot 3 } = \sqrt[3]{162}\] which is not the original problem. so the answer is not A

OpenStudy (phi):

can you do that for the other choices?

OpenStudy (anonymous):

how would i do c like this or \[\sqrt[3]{6*4*4*4*4}\]

OpenStudy (anonymous):

or is that not right

OpenStudy (phi):

4 times itself 3 times (not 4 times) we are using the idea that \[4= \sqrt[3]{4\cdot 4 \cdot 4} \] to "move the 4 inside"

OpenStudy (anonymous):

do i take the 4th 4 away because then it = 384

OpenStudy (anonymous):

wait so is it c then

OpenStudy (phi):

The 4th 4 does not belong there. \[ 4\sqrt[3]{6} = \sqrt[3]{6 \cdot 4 \cdot 4 \cdot 4} \]

OpenStudy (anonymous):

okay so it is c

OpenStudy (phi):

yes

OpenStudy (phi):

the other way to do the problem is factor 384 into its prime factors 2*192= 2*2*96= 2*2*2*48= 2*2*2*2*24 = 2*2*2*2*2*12= 2*2*2*2*2*2*2*3 and pull out triples (2*2*2) * (2*2*2) * 2*3 in other words \[ \sqrt[3]{ 2 \cdot 2 \cdot 2} \sqrt[3]{ 2 \cdot 2 \cdot 2} \sqrt[3]{ 2 \cdot 3} \] which becomes \[ 2 \cdot 2 \sqrt[3]{ 6} \\ 4 \sqrt[3]{ 6}\]

OpenStudy (anonymous):

okay thanks